Stuck on hw. A 50kg firework is shot at 35m/s before exploding, it splits into 2 smaller 25kg pieces. How fast to the smaller pieces move, ignoring air resistance
1
Expert's answer
2020-11-05T10:44:01-0500

The intial momentum of the firework is:


"p_i = mv_i = 50kg\times 35m/s = 1750\space kg\cdot m/s"

According to the momentum conservation law, 2 smaller pieces should have the same momentum in total (in the direction of initial movment):


"p_f = m_1v_1 + m_2v_2 = p_i"

where "m_1 = m_2\equiv m_s = 25kg" is the mass of one piece and "v_1", "v_2" are their velocities. Thus:


"v_1 + v_2 = \dfrac{p_i}{m_s} = \dfrac{1750\space kg\cdot m/s}{25kg} = 70m/s"

If we assume the conservation of kinetic energy (no external forces), then, due to the symmetry will be:


"v_1 = v_2 = \dfrac{70m/s}{2} = 35m/s"

Answer. Both velocities are 35 m/s.